OpenStudy (anonymous):

At a height of h meters, you can see V kilometers to the horizon. These numbers are related by the equation V=3.5√h. A person can see 364 km to the horizon from an airplane window. How high is the airplane? The airplane is ? meters high.

OpenStudy (anonymous):

V^2=3.5^2 * h h= (V/3.5)^2 h=(364/3.5)^2 h=10816